OpenStudy (anonymous):

If the sum of two positive consecutive integers is a, what is the value of the smaller integer in terms of a?

OpenStudy (anonymous):

WHY AND HOW?

OpenStudy (campbell_st):

the mean of the 2 integers is a/2 so to find the lower integer subtract 1/2 from the mean of the integers... so the lower integer is (a -1)/2 here is an example 4 + 5 = 9 9/2 = 4.5 subtract 1/2 leaves the lower integer as 4... hope this make sense

OpenStudy (kropot72):

Another way to get the same result as @campbell_st is as follows: Let the smaller integer = x Then the larger integer = x + 1 x + x + 1 = a 2x = a - 1 x = (a - 1)/2

OpenStudy (anonymous):

You're on the right track @campbell_st, but you showed the problem as 2 consecutive numbers, but the OP stated that the sum is of 2 consecutive POSITIVE integers. So for this you would simply divide the sum by 2, and then taking that quotient (Q), the 2 consecutive positive integers would be Q-1 and Q+1